# Backpropagation-free Spiking Neural Networks with the Forward-Forward Algorithm

### v1

@@ Abstract
Spiking Neural Networks (SNNs) offer a biologically inspired computational paradigm that emulates neuronal activity through discrete spike-based processing. Despite their advantages, training SNNs with traditional backpropagation (BP) remains challenging due to computational inefficiencies and a lack of biological plausibility. This study explores the Forward-Forward (FF) algorithm as an alternative learning framework for SNNs. Unlike backpropagation, which relies on forward and backward passes, the FF algorithm employs two forward passes, enabling localized learning, enhanced computational efficiency, and improved compatibility with neuromorphic hardware. We introduce an FF-based SNN training framework and evaluate its performance across both non-spiking (MNIST, Fashion-MNIST, CIFAR-10) and spiking (Neuro-MNIST, SHD) datasets. Experimental results demonstrate that our model surpasses existing FF-based SNNs by over 5% on MNIST and Fashion-MNIST while achieving accuracy comparable to state-of-the-art backpropagation-trained SNNs. On more complex tasks such as CIFAR-10 and SHD, our approach outperforms other SNN models by up to 6% and remains competitive with leading backpropagation-trained SNNs. These findings highlight the FF algorithm’s potential to advance SNN training methodologies and neuromorphic computing by addressing key limitations of backpropagation.
@@ 1 Introduction
Spiking Neural Networks (SNNs), as a biologically inspired model of neural computation, are designed to process information using discrete spike events [1], closely mimicking the dynamics of real neurons. However, training SNNs remains effectively a significant challenge, primarily due to the credit assignment problem [2] and the difficulty attributing individual neuron contributions to network output errors. Unlike traditional artificial neural networks, SNNs operate on non-differentiable spike dynamics, complicating the application of gradient-based training methods like backpropagation [3]. To address this, surrogate gradient [4] methods are introduced, allowing approximate gradients for effective training while preserving the spike-based behavior of SNNs. The surrogate gradient method replaces the non-differentiable spike activation function with a smooth, differentiable approximation during the backward pass of training. This allows gradients to propagate through the network, enabling optimization using gradient descent.
@@
Despite many successes in training spiking neural networks, backpropagation has faced persistent criticism for its lack of alignment with the mechanisms of the brain [5], raising doubts about its viability as a model for credit assignment in neural systems [6]. One of the major limitations lies in its requirement to explicitly store all neural activity for later use during synaptic adjustments. This explicit storage mechanism has no direct counterpart in biological neural circuits, where memory and computation are more integrated [7]. Another challenge is the need for error derivatives to travel along a dedicated global feedback pathway to create teaching signals. This kind of structured, system-wide error communication contrasts sharply with the distributed and localized nature of information processing in the brain [8]. Furthermore, backpropagation assumes that error signals can move backward through the same neural pathways that forward-propagate information, a phenomenon referred to as the weight transport problem [9]. This assumption is biologically implausible, as neural pathways in the brain are not designed to facilitate such exact bidirectional transport [10]. The sequential nature of backpropagation adds another layer of incompatibility. Unlike the brain’s ability to perform massively parallel computations, backpropagation enforces a process where inference and learning must occur largely step-by-step. This sequentiality significantly diverges from biological neural networks’ highly dynamic and concurrent operations [11]. These challenges highlight why backpropagation is unlikely to serve as a biologically plausible brain learning model. However, recent advancements have been exploring alternative approaches [12] that address these limitations and offer solutions more in line with biological principles.
@@
The Forward-Forward algorithm [13], by inspiration from contrastive methods like Boltzmann Machines [14] and Noise Contrastive Estimation [15], is a novel alternative to the traditional backpropagation method for training neural networks, which Geoffrey Hinton introduced. This algorithm replaces the forward and backward passes of backpropagation with two forward passes. Each layer of the network has its own objective function to optimize the corresponding layer’s weights. FF does not require storing all intermediate activities for backpropagation, nor does it assume symmetrical weight connections. This makes it a more plausible model for cortical learning and suitable for low-power analog hardware. Also, FF can process data sequentially, enabling on-the-fly learning without halting the data pipeline for all the weights’ updating. Another feature of FF is that, unlike backpropagation, it does not require direct connectivity of all consecutive layers through the forward pass, allowing for the inclusion of black-box components within the network.
@@
Motivated by the advantages of the FF algorithm and its potential for SNN training, this work develops an FF-based SNN learning approach and investigates its performance on established datasets. We summarize our contributions as follows:
@@
  • We present a novel learning approach for training spiking neural networks using gradient descent inspired by the Forward-Forward algorithm, which is effective across different types of datasets.
@@
  • We propose a structured training approach for spiking neural networks that leverages key features of the Forward-Forward algorithm, such as layer-level local learning, black-box handling, and more.
@@
  • For the first time, we present the results of an SNN model trained with the FF algorithm on spiking datasets such as N-MNIST and SHD, demonstrating competitive performance against BP-based models and establishing the FF approach as a viable alternative for SNN training.
@@
The rest of this paper is organized as follows: Section 2 overviews related work and backgrounds of SNNs and FF algorithm. Section 3 presents the proposed method. Section 4 elaborates on the experimental setup and results. Finally, Section 5 concludes the paper and suggests a direction for future work.
@@ 2 Related Work and Background
This section presents an overview of relevant literature for this work, including Learning Approaches for SNNs, Forward-Forward Algorithm, and Forward-Forward Algorithm in SNNs.
@@ 2.1 Learning Approaches for SNNs
Since the emergence of spiking neural networks, various ways have been used to train them. These learning approaches are generally divided into three categories: STDP, ANN-SNN conversion, and Backpropagation. STDP [16] is an unsupervised learning algorithm inspired by biological neural systems, which adjusts synaptic weights based on the relative timing of presynaptic and postsynaptic spikes. The fundamental principle is that the order of these spikes determines whether the synapse is strengthened or weakened. In Hebbian STDP, a presynaptic spike occurring before a postsynaptic spike leads to synaptic strengthening, reinforcing the association between neurons. Conversely, in anti-Hebbian STDP, the same spike timing results in synaptic weakening, promoting different forms of neural plasticity [17]. Some variants of the STDP approach include Reward-modulated STDP [18], Mirrored STDP [19], and Probabilistic STDP [20], each of which can be utilized for different purposes. ANN-to-SNN conversion approach is a widely used method for leveraging the strengths of both networks [21]. It involves training a standard ANN using traditional methods and then adapting it into an SNN by mapping its weights and neuron parameters to spiking equivalents. This approach combines the efficiency of ANN training with the benefits of SNN operation, allowing SNNs to achieve performance comparable to ANNs on some benchmarks.
@@ 2.2 Forward-Forward Algorithm
The Forward-Forward Algorithm is a novel learning approach for neural networks that forgoes traditional backpropagation by using two forward passes: one with real or ”positive” data and another with ”negative” or contrasting data. First, the algorithm prepares data by creating positive samples, which are real labeled data, and generating negative samples. These negative samples are carefully crafted, either by corrupting positive samples or using synthetic data, to differ structurally from positive data. In supervised learning, in order to use labels in the training process, positive and negative samples can be generated from the original data in various ways. One common way is to embed the corresponding labels on each sample. When embedding, it should be noted that if we are creating positive samples, we must put the correct label on the data, and if we are looking to create negative samples, we must put the wrong label on the data. Fig. 1 shows how to embed labels and create positive and negative samples corresponding to a sample from class one from a dataset that has 10 classes. It can be seen that in the positive sample corresponding to the selected sample from class one, the second pixel is lit out of the first ten pixels. Similarly, in general, in creating its corresponding negative sample, any pixel except the second pixel can be lit out of the first ten pixels, which, for example, class zero is considered as the class of the negative sample.
@@ 2.3 Forward-Forward Algorithm in SNNs
Despite the widespread use of the BP method in learning spiking neural networks, problems such as weight transport, update locking, and black-box handling still persist. Using the Forward-Forward algorithm instead of backpropagation and replacing its steps in the training process of spiking neural networks can solve these problems to a good extent. Studies have also been conducted using the FF algorithm’s learning idea to spiking neural networks. Ororbia [27] introduces a novel neuro-mimetic architecture of SNNs designed to operate with layer-wise parallelism and synaptic plasticity. The proposed model, termed contrastive-signal-dependent plasticity (CSDP), generalizes forward-forward learning for spiking systems by iteratively processing sensory inputs within dynamic, recurrent circuits. This system uniquely avoids feedback pathways, offering biologically inspired and computationally efficient mechanisms for classification and reconstruction tasks. The architecture integrates a layer-wise parallel design incorporating bottom-up, top-down, and lateral signals for dynamic computation without forward-locking and introduces CSDP as an online, biologically plausible plasticity mechanism for adapting synapses using local spike-based activity traces. Terres et al. [28] explore using the Forward-Forward algorithm as a biologically plausible alternative to backpropagation, specifically adapting it for SNNs to enhance energy efficiency and robustness. The authors present two spiking-specific goodness functions—bounded and unbounded—tailored to the discrete spiking activity of neurons, enabling the effective integration of the FF algorithm into SNNs. Also, building on these properties, they introduce the FF-SCP algorithm, a novel approach for out-of-distribution detection.
@@ 3 Proposed Method
The training procedure for SNNs using the forward-forward algorithm involves several crucial steps to ensure effective learning and optimal performance. To create training examples, we superimpose the labels onto the input images. This process involves setting the first $n$ elements of the flattened input vector to zero which $n$ is the number of classes. Then assign the maximum pixel value of the image to the index corresponding to the true label. This labeling aids in distinguishing between different digit classes during the training. We use hard labeling to assign the labels of negative samples. For each input image, the model computes a ”goodness” score for all possible class labels. The goodness score measures the mean squared activation of the network layers in response to the input. The goodness score for the true class label of each input image is set to zero. This ensures that the true label is not considered when selecting a hard label. The remaining goodness scores (excluding the true label) are transformed using a square root function. This transformation flattens the distribution, making it less peaked and more uniform, thus avoiding the dominance of any single class label. Using the transformed goodness scores as a probability distribution, a hard label is sampled for each input image. This label is chosen from the set of non-true labels and represents a class that the network finds relatively difficult to distinguish from the true label. Hard labeling aims to create challenging negative examples for the network.
After generating positive and negative samples, it’s time to feed them into the network. The network comprises an input layer and hidden layers that are fully connected. Notably, there is no output layer, but the number of hidden layers can be increased according to the problem’s needs. For the non-spiking datasets, the training data vectors are directly encoded into the first layer, whereas for the spiking datasets, the data is entered with their respective coding. After a linear transformation of the input data, the desired neuron model, like leaky integrate-and-fire (LIF), is performed. In the Forward-Forward algorithm, the primary objective is to enhance neuronal activation in response to positive data while reducing activation in response to negative data. However, if the output of the neurons is directly propagated to the subsequent hidden layer, it may hinder effective learning in deeper layers, particularly when processing positive data. This issue arises due to the influence of large values carried over from the preceding layer. To mitigate this effect and ensure that deeper hidden layers learn independently of prior layers, a normalization step is applied before activating the mechanism in the Leaky Integrate-and-Fire model. LIF neurons accumulate input current over time and emit a spike when the membrane potential exceeds a predefined threshold. The membrane potential then resets, mimicking the functions observed in biological neurons. The neurons are initialized with parameters like leakage term and spiking threshold. The recursive representation of a LIF neuron is as follows:
@@
where $U$, $W$, $X$, $R$, and $\beta$ are membrane potential, weights, inputs, reset term, and decay rate, respectively. In each time step, spike emission occurs if the membrane potential exceeds the membrane potential threshold $U_{thr}$. The following equation represents this:
$S[t]=\begin{cases}1&U_{thr}\leq U[t],\\ 0&otherwise.\end{cases}$ (2)
Our method has two forward passes called positive and negative passes. In the positive pass, positive data are entered into the network, and for each step, the condition of the occurrence of spikes is determined. We calculate the goodness value of each hidden layer for positive data by considering the spikes’ count of each neuron as the input of the goodness function. After that, the negative pass starts, and negative samples are entered into the neural network to calculate the goodness of the negative samples by going through the same steps as the positive data. During the execution of positive and negative forward passes, for each neuron, the spike count value is equal to the sum of all spikes through all time steps and calculated as follows:
$C_{pos}=\sum\limits_{t=1}^{T}S_{pos}[t],\$ (3)
$C_{neg}=\sum\limits_{t=1}^{T}S_{neg}[t],\$ (4)
where $T$ is the number of time steps. Also, $S_{pos}$ and $S_{neg}$ are spike occurrences of neurons in the face of positive and negative samples, respectively. In each layer, the goodness function $G$ for positive and negative samples is defined as follows:
$G_{pos}=\frac{1}{N}\sum\limits_{i=1}^{N}C_{pos,i}^{2},\$ (5)
$G_{neg}=\frac{1}{N}\sum\limits_{i=1}^{N}C_{neg,i}^{2},\$ (6)
where N is the number of neurons in the desired layer, and $C_{pos,i}$ and $C_{neg,i}$ are the spike counts of neuron $i$ for positive and negative data, respectively. Using the goodness values obtained from positive and negative samples, the loss of the layer is calculated with a formula which is inspired by the swish function [29] and described as follows:
$\Delta=G_{\text{pos}}-G_{\text{neg}},$ (7)
@@
where $\alpha$ is a scaling factor. Training of spiking neural networks is challenging due to the non-differentiability of the spike generation process. To overcome this, surrogate gradient learning can be utilized. Surrogate gradients provide a differentiable approximation of the spike function, allowing gradient-based optimization methods to be used. During layer-wise weight updating, these surrogate gradients approximate the true gradients of the spike function, enabling the network to update its parameters effectively. Various functions can be considered for gradient approximation, the most important of which are sigmoid, fast sigmoid, and shifted arc-tan. Refer to equation (2), shifted arc-tangent surrogate function and its gradient w.r.t. the membrane potential is defined as follows:
$S$ $\approx\frac{1}{\pi}arctan(\pi U\frac{\alpha}{2})$ (8)
$\frac{\partial S}{\partial U}$ $=\frac{1}{\pi}\frac{1}{(1+(\pi U\frac{\alpha}{2})^{2})}$
where $U$ is the membrane potential and $k$ is the slope value, a hyperparameter. Obviously, other proposed approximation functions can also be used to calculate the surrogate gradient. After updating the weights of each layer, the next epoch of training is implemented.
@@
Fig. 2 illustrates the learning procedure of the proposed method. It is assumed that the training operation is to be performed on an original sample. By applying the label embedding operation, the sample is converted to a positive and negative sample, and before entering the fully connected neural network, each positive and negative sample must be flattened and converted to a vector. In the positive forward pass, the positive sample enters the neural network, and spikes are generated in the neurons during the time steps. In each neuron, the number of spikes is counted independently, and after that, their spike count passes to goodness function $G$ to calculate the goodness value of the positive sample. For the negative sample, the same steps are followed, and at the end, the goodness of the negative sample is calculated. After finding the goodness values for positive and negative samples, the amount of loss is calculated with the loss function $L$, and then the weights of the corresponding layer are updated locally.

### v2

@@ Abstract
Spiking Neural Networks (SNNs) offer a biologically inspired computational paradigm that emulates neuronal activity through discrete spike-based processing. Despite their advantages, training SNNs with traditional backpropagation (BP) remains challenging due to computational inefficiencies and a lack of biological plausibility. This study explores the Forward-Forward (FF) algorithm as an alternative learning framework for SNNs. Unlike backpropagation, which relies on forward and backward passes, the FF algorithm employs two forward passes, enabling layer-wise localized learning, enhanced computational efficiency, and improved compatibility with neuromorphic hardware. We introduce an FF-based SNN training framework and evaluate its performance across both non-spiking (MNIST, Fashion-MNIST, Kuzushiji-MNIST) and spiking (Neuro-MNIST, SHD) datasets. Experimental results demonstrate that our model surpasses existing FF-based SNNs on evaluated static datasets with a much lighter architecture while achieving accuracy comparable to state-of-the-art backpropagation-trained SNNs. On more complex spiking tasks such as SHD, our approach outperforms other SNN models and remains competitive with leading backpropagation-trained SNNs. These findings highlight the FF algorithm’s potential to advance SNN training methodologies by addressing some key limitations of backpropagation.
@@ 1 Introduction
Spiking Neural Networks (SNNs), as a biologically inspired model of neural computation, are designed to process information using discrete spike events [1], closely mimicking the dynamics of real neurons. However, training SNNs remains a significant challenge, primarily due to the credit assignment problem [2] and the difficulty of attributing the contributions of individual neurons to network output errors. Unlike traditional artificial neural networks, SNNs operate on non-differentiable spiking activities, complicating the application of gradient-based training methods like backpropagation [3]. To address this, surrogate gradient [4] methods are introduced, allowing approximate gradients for effective training while preserving the spike-based behavior of SNNs. The surrogate gradient method replaces the non-differentiable spike activation function with a smooth, differentiable approximation during the backward pass of training. This enables the use of gradient-based learning algorithms such as backpropagation in the training process of SNNs.
@@
Despite notable successes in training spiking neural networks, backpropagation has faced persistent criticism for its training challenges and lack of alignment with the mechanisms of the brain [5], raising doubts about its viability as a model for credit assignment in neural systems [6]. One of the major limitations lies in its requirement to explicitly store all neural activity for later use during synaptic adjustments. This explicit storage mechanism has no direct counterpart in biological neural circuits, where memory and computation are more integrated [7]. Another significant challenge is the requirement for error derivatives to propagate through a dedicated global feedback pathway within the network to generate teaching signals. This kind of structured, system-wide error communication contrasts sharply with the distributed and localized nature of information processing in the brain [8]. It may, in certain cases, exacerbate issues such as vanishing and exploding gradients in algorithmic implementations. Furthermore, backpropagation assumes that global error signals can move backward through the same neural pathways used for forward propagation, a phenomenon referred to as the weight transport problem [9]. This assumption is biologically implausible, as neural pathways in the brain are not designed to facilitate such exact bidirectional transport [10]. The sequential nature of backpropagation adds another layer of incompatibility. Unlike the brain’s ability to perform massively parallel computations, backpropagation enforces a process where inference and learning must occur largely step-by-step. This sequentiality significantly diverges from biological neural networks’ highly dynamic and concurrent operations [11]. These challenges highlight why backpropagation is unlikely to serve as a biologically plausible brain learning model. However, recent advancements have been exploring alternative approaches [12] that address some of these limitations and offer solutions more in line with biological principles.
@@
The Forward-Forward algorithm [13], inspired by contrastive methods like Boltzmann Machines [14] and Noise Contrastive Estimation [15], is a novel alternative to the traditional backpropagation method for training neural networks, which Geoffrey Hinton introduced. This algorithm replaces backpropagation’s forward and backward passes with two forward passes. Each network layer has its own objective function to optimize the corresponding layer’s weights. FF does not require storing all intermediate activities for updating each layer’s weights, nor does it assume symmetrical weight connections. This makes it a more plausible model for cortical learning and suitable for low-power analog hardware [13]. Also, FF can process data sequentially, enabling on-the-fly learning without halting the data pipeline for all the weights’ updating. Another feature of FF is that, unlike backpropagation, it does not require direct connectivity of all consecutive layers through the forward pass, allowing for the inclusion of black-box components within the network.
@@
Motivated by the advantages of the FF algorithm and its potential for SNN training, this work develops an FF-based SNN learning approach and investigates its performance on static and spiking benchmark datasets. We summarize our contributions as follows:
@@
  • A novel gradient-based learning approach is introduced for training spiking neural networks inspired by the Forward-Forward algorithm, which demonstrates strong performance across a range of benchmark datasets.
@@
  • A structured training approach for spiking neural networks that leverages key principles of the Forward-Forward algorithm, including layer-wise local learning, black-box compatibility, and additional advantageous features.
@@
  • Experimental results of our SNN model, trained using the FF algorithm on spiking datasets such as N-MNIST and SHD, demonstrate competitive performance compared to backpropagation-based models and establish the FF approach as a viable alternative for training spiking neural networks.
@@
The rest of this paper is organized as follows: Section 2 overviews related work and backgrounds of SNNs and the FF algorithm. Section 3 presents the proposed method. Section 4 elaborates on the experimental setup and results. Finally, Section 5 concludes the paper and suggests a direction for future work.
@@ 2 Related Work and Background
This section provides an overview of the literature relevant to our work, covering learning approaches for spiking neural networks, the Forward-Forward algorithm, and its application within the context of SNNs.
@@ 2.1 Learning Approaches for SNNs
Since the emergence of spiking neural networks, various ways have been used to train them. These learning approaches are generally divided into three categories: STDP, ANN-SNN conversion, and Backpropagation. STDP [16] is an unsupervised learning algorithm inspired by biological neural systems, which adjusts synaptic weights based on the relative timing of presynaptic and postsynaptic spikes. The fundamental principle is that the order of these spikes determines whether the synapse is strengthened or weakened. In Hebbian STDP, a presynaptic spike occurring before a postsynaptic spike leads to synaptic strengthening, reinforcing the association between neurons. Conversely, in anti-Hebbian STDP, the same spike timing results in synaptic weakening, promoting different forms of neural plasticity [17]. Some variants of the STDP approach include Reward-modulated STDP [18], Mirrored STDP [19], and Probabilistic STDP [20], each of which can be utilized for different purposes.
ANN-to-SNN conversion approach is a widely used method for leveraging the strengths of both networks [21]. It involves training a standard ANN using traditional methods and then adapting it into an SNN by mapping its weights and neuron parameters to spiking equivalents. This approach combines the efficiency of ANN training with the benefits of SNN operation, allowing SNNs to achieve performance comparable to ANNs on some benchmarks.
@@ 2.2 Forward-Forward Algorithm
The Forward-Forward Algorithm is a novel learning approach for neural networks that forgoes traditional backpropagation by using two forward passes: one with real or ”positive” data and another with ”negative” or contrasting data. First, the algorithm prepares data by creating positive samples, which are real labeled data, and generating negative samples. These negative samples are carefully crafted, either by corrupting positive samples or using synthetic data, to differ structurally from positive data. In supervised learning, in order to use labels in the training process, positive and negative samples can be generated from the original data in various ways. One common way is to embed the corresponding labels on each sample. When embedding, it should be noted that if we are creating positive samples, we must put the correct label on the data, and if we are looking to create negative samples, we must put the wrong label on the data.
Figure 1 shows how to embed labels into an image from a static dataset and create positive and negative samples corresponding to a sample from class one from a dataset that has 10 classes. It can be seen that in the positive sample corresponding to the selected sample from class one, the second pixel is lit out of the first ten pixels. Similarly, in general, in creating its corresponding negative sample, any pixel except the second pixel can be lit out of the first ten pixels, which, for example, class zero is considered as the class of the negative sample.
@@ 2.3 Forward-Forward Algorithm in SNNs
Despite the widespread use of the BP method in learning spiking neural networks, problems such as weight transport, update locking, and black-box handling still persist. Using the Forward-Forward algorithm instead of backpropagation and replacing its steps in the training process of spiking neural networks can solve these problems to a good extent. Studies have also been conducted using the FF algorithm’s learning idea to spiking neural networks.
Ororbia [27] propose a biologically plausible learning framework for spiking neural networks called Contrastive Signal–Dependent Plasticity (CSDP). This method enables self-supervised learning by combining local plasticity rules with contrastive principles derived from forward-forward learning. The architecture consists of recurrent layers of leaky integrate-and-fire neurons that are connected via top-down, bottom-up, and lateral synapses. These layers operate in parallel and asynchronously, allowing for efficient and scalable simulation while avoiding biologically implausible constraints such as feedback connections or backpropagation through time. At the core of the learning process is an activation trace that smooths sparse spike activity over time. Each layer adapts its synapses based on a local contrastive objective that encourages the network to increase the “goodness” or likelihood of in-distribution (positive) inputs and suppress the response to out-of-distribution (negative) inputs. This is achieved using a Hebbian-like synaptic update rule modulated by a local contrastive signal, eliminating the need for non-local information such as global error signals or explicit error feedback pathways. Negative samples required for contrastive learning are generated dynamically. In the supervised setting, they are created by randomly assigning incorrect labels to cloned inputs. In the unsupervised case, they are synthesized using random augmentations and interpolations of input patterns within a batch. The model is further extended to perform task-specific functions such as classification and reconstruction by adding simple generative and predictive synaptic pathways that follow similar local learning rules.
Terres et al. [28] introduces a framework for robust Out-of-Distribution (OoD) detection using fully-spiking neural networks trained with the Forward-Forward algorithm, a forward-only alternative to backpropagation. They adapt FF to the spiking domain by proposing two spiking-compatible goodness functions. One is unbounded to mimic ReLU behavior via spike aggregation, and one is bounded, aligning with the discrete, time-limited spiking activity. They then propose FF-SCP, a new distance-based OoD detection algorithm that uses the latent space structure induced by FF. Instead of relying on class centroids, FF-SCP calculates the minimum distance between a sample’s latent representation and class-specific latent manifolds, providing a more accurate reflection of class boundaries. This distance serves as an OoD score. To improve explainability, the authors further design a gradient-free attribution technique that uses a decoder to reconstruct inputs from latent vectors. By optimizing these latent representations to match a target class distribution while staying close to the original input, the method generates attribution maps that highlight features responsible for OoD classification. Experiments across datasets such as MNIST, F-MNIST, K-MNIST, and Not-MNIST show that the proposed spiking FF system achieves OoD detection performance comparable to or better than state-of-the-art spiking and non-spiking methods like SCP and ODIN, with strong energy-efficiency implications due to the spiking architecture.
@@ 3 Proposed Method
The spiking neural networks training procedure using the Forward-Forward algorithm involves generating labeled training samples that align with the algorithm’s layer-wise contrastive learning framework. A critical step in this process is superimposing class labels onto the input vectors by replacing part of the original input with class-specific information. This enables the network to evaluate whether a given input-label pair is compatible, allowing each layer to independently distinguish between positive (correct) and negative (incorrect) associations. Consider a flattened input vector $x^{input}\in\mathbb{R}^{d}$ like:
$x^{input}=[x_{0},x_{1},\dots,x_{d-1}].$ (1)
To create these labeled inputs from both static and spatiotemporal data, we modify the input vector by replacing the first $c$ elements of the flattened vector, where $c$ is the number of classes, with a label-encoding vector. Specifically, we initialize these $c$ elements to zero and assign the maximum element’s value of the original input, denoted as $m$, to the index corresponding to the label class:
$m=\max_{0\leq i\leq d-1}x^{\text{input}}_{i},$ (2)
the remaining $d-c$ elements retain the corresponding values from the original input vector $x^{input}$. The modified vector ${x^{pos}}\in\mathbb{R}^{d}$ for the positive sample with true label $y^{pos}=t$ is constructed as:
$x_{i}^{pos}=\begin{cases}0&i\leq c\ and\ i\neq t,\\ m&i=t,\\ x_{i}&i\geq c,\end{cases}$ (3)
where the value $m$ replaces the $t^{th}$ entry (index $t-1$) to indicate the true label, while the remaining elements $x_{c}$ to $x_{d-1}$ are directly copied from the original input. To construct a negative sample, we choose a false label $y^{\text{neg}}=f\neq t$ , and embed it as:
$x_{i}^{neg}=\begin{cases}0&i\leq c\ and\ i\neq f,\\ m&i=f,\\ x_{i}&i\geq c.\end{cases}$ (4)
@@
In this case, the embedded label does not match the ground truth, and the network is expected to learn to distinguish positive samples from negative samples. This label replacement strategy embeds class identity directly into the input vector without changing its dimensionality. It supports forward-only contrastive training by enabling each layer to compute a scalar goodness score and adjust weights locally, eliminating the need for explicit output layers or backpropagated error signals.
To choose the index $f$ for negative samples more intelligently instead of randomly, we use hard labeling to assign the labels of negative samples. For each input image, the model computes a ”goodness” score for all possible class labels. The goodness score measures the mean squared activation of the network layers in response to the input. The goodness score for the true class label of each input image is set to zero. This ensures that the true label is not considered when selecting a hard label. The remaining goodness scores (excluding the true label) are transformed using a square root function. This transformation flattens the distribution, making it less peaked and more uniform, thus avoiding the dominance of any single class label. Using the transformed goodness scores as a probability distribution, a hard label is sampled for each input image. This label is chosen from the set of non-true labels and represents a class that the network finds relatively difficult to distinguish from the true label. Hard labeling aims to create challenging negative examples for the network. After generating the positive and negative samples, they are fed into the network, which consists of an input layer followed by fully connected hidden layers. Notably, the network does not include an output layer, and the number of hidden layers can be adjusted based on the complexity of the task. Following a linear transformation of the input data, a specified neuron model—such as the leaky integrate-and-fire (LIF) model—is applied. LIF neurons accumulate input current over time and emit a spike when the membrane potential exceeds a predefined threshold. The membrane potential then resets, mimicking the functions observed in biological neurons. The neurons are initialized with parameters like the leakage term and the spiking threshold. The recursive representation of a LIF neuron is as follows:
@@
where $U$, $W$, $X$, $R$, and $\beta$ are membrane potential, weights, inputs, reset term, and decay rate, respectively. In each time step, spike emission occurs if the membrane potential exceeds the threshold $U^{thr}$:
$S[t]=\begin{cases}1&U^{thr}\leq U[t],\\ 0&otherwise.\end{cases}$ (6)
The Forward-Forward algorithm aims to increase neuronal activation in response to positive samples while suppressing activation for negative samples in every layer. However, directly propagating neuron outputs to the next hidden layer can impede effective learning in deeper layers, especially for positive samples. This challenge stems from the accumulation of large activations from preceding layers, which can dominate the learning signal. To address this, a normalization step [29] is applied to the output of each layer’s linear transformation (i.e., the input to the subsequent layer) at every time step, allowing each hidden layer to learn independently of its predecessors. In each layer, the normalization performed at every time step $t$ is defined as follows:
${N}(WX[t])=\gamma[t]\odot\frac{WX[t]-\mu[t]}{\sqrt{\sigma^{2}[t]+\epsilon}}+\beta[t],$ (7)
where the parameters $\gamma$ and $\beta$ are learned during training. The constant $\epsilon$ is a small value added to the variance to prevent division by zero during normalization, and the symbol $\odot$ represents the Hadamard product, an element-wise multiplication operation. $\mu$ and $\sigma^{2}$ are the mean and variance along the samples of the mini-batch, respectively, and for each layer, they are defined as follows:
$\mu[t]$ $=\frac{1}{B}\sum_{b=1}^{B}WX_{b}[t],$ (8)
$\sigma^{2}[t]$ $=\frac{1}{B}\sum_{b=1}^{B}\left(WX_{b}[t]-\mu[t]\right)^{2},$
where $B$ is the batch size. Accordingly, the formulation of the LIF neuron model can be reformulated as:
$U[t+1]=\beta U[t]+N(WX[t+1])-R[t].$ (9)
Our method involves two distinct forward passes, the positive and negative passes. During the positive pass, positive samples are presented to the network, and at each time step, the spiking condition for each neuron is evaluated. The goodness value for each hidden layer is then computed by applying a goodness function to the spike count of each neuron. Subsequently, the negative pass is performed, where negative samples are propagated through the network using the same procedure to compute their corresponding goodness values. Throughout both the positive and negative forward passes, the spike count for each neuron $i$, defined as the total number of spikes accumulated over all time steps, is computed as follows:
$C_{i}^{pos}=\sum\limits_{t=1}^{T}S_{i}^{pos}[t],\$ (10)
$C_{i}^{neg}=\sum\limits_{t=1}^{T}S_{i}^{neg}[t],$ (11)
where $T$ denotes the total number of time steps, while $S_{i}^{\text{pos}}$ and $S_{i}^{\text{neg}}$ represent the spike occurrences of neuron $i$ in response to positive and negative samples, respectively. In each layer, the goodness function $G$ for positive and negative samples is defined as follows:
$G^{pos}=\frac{1}{N}\sum\limits_{i=1}^{N}(C^{pos}_{i})^{2},\$ (12)
$G^{neg}=\frac{1}{N}\sum\limits_{i=1}^{N}(C^{neg}_{i})^{2},\$ (13)
where $N$ is the number of neurons in the given layer. Based on the goodness values computed for positive and negative samples, the loss for each layer is calculated independently using a function inspired by the Swish function [30], defined as follows:
$\Delta=G^{\text{pos}}-G^{\text{neg}},$ (14)
@@
Figure 3: Loss function $\frac{-\alpha\Delta}{1+e^{\alpha\Delta}}$ plotted against $\Delta$ for different scaling factor values of $\alpha$.
where $\alpha$ is a scaling factor that controls the sharpness of the loss curve. Figure 3 illustrates the behavior of the loss function plotted against the variable $\Delta$ for different values of the scaling factor $\alpha\in\{1,2,5\}$. This function serves as a smooth contrastive objective, designed to increase the margin between positive and negative goodness values. As depicted, for negative values of $\Delta$, the loss increases approximately linearly, with steeper slopes for larger values of $\alpha$. This implies stronger penalization when the goodness of negative samples surpasses that of positive samples. In contrast, for the large positive values of $\Delta$, the loss asymptotically approaches zero, indicating that well-separated positive and negative samples contribute less to the gradient, thereby preventing excessive updates once correct classification with sufficient margin has been achieved. Higher values of $\alpha$ result in sharper transitions near $\Delta=0$, effectively tightening the decision boundary and making the loss function more sensitive to misclassified or marginal examples. This makes $\alpha$ a critical hyperparameter for modulating margin sensitivity and learning dynamics. The weights of each layer are updated to minimize the loss value by computing the gradient of the loss with respect to the layer’s weights. Training of spiking neural networks is challenging due to the non-differentiability of the spike generation process. To overcome this, surrogate gradient learning [4] can be utilized. Surrogate gradients provide a differentiable approximation of the spike function, allowing gradient-based optimization methods to be used. During layer-wise weight updating, these surrogate gradients approximate the true gradients of the spike function, enabling the network to update its parameters effectively. Guided by experimental results, we utilize the shifted arctangent surrogate function [21], with its gradient with respect to the membrane potential expressed as:
$S[t]$ $\approx\frac{1}{\pi}arctan(\pi\frac{\alpha}{2})U[t],$ (15)
$\frac{\partial S[t]}{\partial U[t]}$ $=\frac{1}{\pi}\frac{1}{(1+(\pi\frac{\alpha}{2}U[t])^{2})},$
@@
where $\alpha$ is the slope parameter. After updating the weights of each layer, the next epoch of training is implemented. Figure 2 illustrates the learning procedure of the proposed method. It is assumed that the training operation is to be performed on an original sample. By applying the label embedding operation, the sample is converted to a positive and negative sample, and before entering the fully connected neural network, each positive and negative sample must be flattened and converted to a vector. In the positive forward pass, the positive sample enters the neural network, and spikes are generated in the neurons during the time steps. In each neuron, the number of spikes is counted independently, and after that, their spike count passes to the goodness function $G$ to calculate the goodness value of the positive sample. For the negative sample, the same steps are followed, and at the end, the goodness of the negative sample is calculated. After finding the goodness values for positive and negative samples, the amount of loss is calculated with the loss function $L$, and then the weights of the corresponding layer are updated locally.
During inference, the absence of a dedicated output layer precludes the use of conventional classification methods. To address this, we adopt a label-scoring strategy that leverages layer-wise information for prediction. Specifically, we iterate over all possible class labels for each test sample by embedding each label into the input, as done during training. Each resulting input-label pair is passed through the trained network, and the goodness values computed across all layers are summed. The label that yields the highest total goodness score is then selected as the predicted class for the sample.